Project

General

Profile

Physics Shift Report

Start Date: 23rd February
End Date: 27th February
Start Run: 07620
End Run: 07652
Physics shifter: Francois Drielsma

Summary

23/02: First day of data taking was divided in several activities
  • Beam bump tuning with H.N. and D. Adams in the morning
  • Shifter training from 2PM to 3:30PM
  • TOF calibration data taking until 10PM (no DS, data looks fine from onRec), reached 500k TOF1 trig.
Run number Run type Mom. setting Trigger condition Nbr. of spills Int. part. tr.
7623 Pion + 200 MeV/c TOF1 7133 175664
7624 Pion + 200 MeV/c TOF1 6140 166314
7625 Pion + 200 MeV/c TOF1 6316 161340
24/02: Second day of data taking
  • FC being bolted to the SSs, evacuation of the absorber cavity (all day)
  • /!\ No reference run taken !
  • Resume data taking operations, alignment data, reached 550k out of 1M
Run number Run type Mom. setting Trigger condition Nbr. of spills Int. part. tr.
7626 Pion + 300 MeV/c TOF1 3626 166015
7627 Pion + 300 MeV/c TOF1 3669 168548
7628 Pion + 300 MeV/c TOF1 6273 249089
25/02: Third day of data taking
  • Work in the hall during the day to evacuate the FC cavity
  • Resume alignment data taking at 7PM, reached 750k integrated total particle triggers for the first setting, 23k for the second setting.
  • /!\ Proton absorber too thin for all the runs that day (29 mm instead of 82 mm), must be set manually. Might be some proton contamination in the beam up to TOF1 at most (5 cm of PVT in TOF0).
Run number Run type Mom. setting Trigger condition Nbr. of spills Int. part. tr.
7635 Pion + 300 MeV/c TOF1 3716 224796
7636 Muon + 200 MeV/c TOF1 5763 22728
26/02: Fourth day of data taking
  • Work in the hall during the day for the FC and setting up the DS power supply
  • Resume alignment data taking at 4PM, reach 50k for the second setting (very poor rate without DS), 150k for the third setting.
Run number Run type Mom. setting Trigger condition Nbr. of spills Int. part. tr.
7639 Muon + 200 MeV/c TOF1 6014 27685
7640 Pion + 450 MeV/c TOF1 3266 40194
7641 Pion + 450 MeV/c TOF1 8889 111295
27/02: Fifth day of data taking
  • Resume alignment data taking at 4PM, reached 150k for the second setting and 250k for the third setting (complete).
    */!\ Wrong tag for 7649, it is 450 pions and not 6pi-240
Run number Run type Mom. setting Trigger condition Nbr. of spills Int. part. tr.
7646 Pion + 450 MeV/c TOF1 2524 24253
7647 Muon + 200 MeV/c TOF1 1802 8003
7648 Muon + 200 MeV/c TOF1 4462 20441
7649 Pion + 450 MeV/c TOF1 2567 35197
7650 Muon + 200 MeV/c TOF1 13566 61760
7651 Pion + 450 MeV/c TOF1 2477 34223
7652 Muon + 200 MeV/c TOF1 5395 27878

Grand total of alignment data collected:

Run type Mom. setting Trigger condition Nbr. requested Int. part. tr.
Pion + 300 MeV/c TOF1 1000k 750k
Muon + 200 MeV/c TOF1 250k 150k
Pion + 450 MeV/c TOF1 250k 250k
Muon + 350 MeV/c TOF1 250k 0k

Run documentation

Elog

  • Entries for each shift and major issues/faults. No problems found, very thorough.
  • Check list filled on a regular basis, good practice

Run Condition Summary Spreadsheet/CDB

  • All of the information recorded in the Summary Spreasheet
  • Run 7621 (calibration run) missing in the CDB

Analysis

  • No particular analysis requested by the measurement coordinator
  • The alignment analysis needs a resurvey of the downstream PID detector to take place.
  • Amount of statistics collected for the TOF calibration from the 7623-7625:


Reconstruction Quality

  • The physics devil processed the files successfully and reported a reasonable number of unpacking errors (< 0.1%)
  • The files have been properly uploaded to the grid
  • The on line reconstruction plots don't reveal any DAQ issues, albeit a noticeable inefficiency in tracker 1 station 1. Might be solved by updating the cabling file, TBD.

imageSciFi_SpacePointsTwoD_tracker_1_station_1_deadarea.png (20.6 KB) Drielsma, François, 25 February 2016 10:07

imageSciFi_Spacepoints_7628.png (11.3 KB) Drielsma, François, 25 February 2016 10:07

EMRReduce_occ_yz_7628.png (52.9 KB) Drielsma, François, 25 February 2016 10:07

EMRReduce_beam_profile_7628.png (32.3 KB) Drielsma, François, 25 February 2016 10:07

CkovReduce_charge_7628.png (17.4 KB) Drielsma, François, 25 February 2016 10:07

TofReduce_tof0_spxy_7628.png (43.8 KB) Drielsma, François, 25 February 2016 10:07

TofReduce_tof1_spxy_7628.png (39.3 KB) Drielsma, François, 25 February 2016 10:07

TofReduce_tof2_spxy_7628.png (51.4 KB) Drielsma, François, 25 February 2016 10:07

KLReduce_7628.png (11.5 KB) Drielsma, François, 25 February 2016 10:07

imageSciFi_SpacePointsTwoD_tracker_0_station_1_7628.png (20.7 KB) Drielsma, François, 25 February 2016 10:07

tof0_sps_small.png (76.9 KB) Drielsma, François, 25 February 2016 15:49

tof1_sps_small.png (63.7 KB) Drielsma, François, 25 February 2016 15:49

tof2_sps_small.png (74.9 KB) Drielsma, François, 25 February 2016 15:49